Maya is 2 km offshore in a boat and wishes to reach a coastal village which is 6 km down a straight
shoreline from the point on the shore nearest to the boat. She can row at 2 km/hr and run at 5
km/hr. Where should she land her boat to reach the village in the least amount of time?
1
Expert's answer
2019-11-13T11:48:06-0500

Finding the Unknown value in the word problem


We are going to find where, Maya land her boat to reach the village in the least amount of time


Solution:


Let C = 2 km and b = 6 km


Speed at running = 5 kmph


speed to walk = 2kmph





By Pythagorean fomrula


"l^2 = c^2 + x^2"

"l = \sqrt {4+x^2}"


Time required to cross the river

"Time = \frac {Distance } { speed} = \frac {l} {2} = \frac {\sqrt {4 +x^2}} {2}"

Maya should run the distance =

"b - x = 6 - x"

Time for run =

"\frac {Distance} {speed} = \frac {6-x}{5}"

Total time =

"t = \frac {1}{2} \times \sqrt {x^2+4} + \frac {6-x} {5}"


"t = \frac {1}{2} \times {(x^2+4)}^{\frac {1}{2}} + \frac {6-x} {5}"

Differentiate with respect to x,

"\frac {dt}{dx} = \frac{1}{2} \times \frac {1}{2} \times {(x^2+4)}^{\frac {1}{2}-1} (2x) - \frac {1}{5}"


"\frac {dt}{dx} = \frac {x}{2} \times {(x^2+4)}^{\frac {1}{2}-1} - \frac {1}{5}"

"\frac {dt}{dx} = \frac {x}{2} \times {(x^2+4)}^{\frac {-1}{2}} - \frac {1}{5}"

"\frac {dt}{dx} = \frac {x}{2} \times \frac {1} {\sqrt {(x^2+4)}} - \frac {1}{5}"

Now set the derivative to zero, to find the minimum value of x



"\frac {x}{2} \times \frac {1} {\sqrt {(x^2+4)}} - \frac {1}{5} =0"

"\frac {x} {\sqrt {(x^2+4)}} = \frac {2}{5}"

Squaring on both the sides,


"\frac {x^2} {{(x^2+4)}} = \frac {4}{25}"

"25x^2 = 4 (x^2 + 4)"

"25x^2 = 4x^2 + 16""21x^2 = 16"

"x^2 = \frac {16} {21}"

"x = \frac {4} {\sqrt 21}"

Answer:


At "x = \frac {4} {\sqrt 21}" km down the shore she land her boat to reach the village.